OpenStudy (anonymous):

A pancake recipe requires three and one half cups of milk to one cup of flour. If four and two thirds cups of milk is used, what quantity of flour will be needed, according to the recipe?

OpenStudy (anonymous):

M2/M1 = F2/F1 (4 2/3) / (3 1/2) = F2 / 1 F2 = (4 2/3) / (3 1/2) F2 =?

OpenStudy (anonymous):

16/9,

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

F2 = 1 1/3 cups of flour

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob